# InterDyn: Controllable Interactive Dynamics with Video Diffusion Models

### v1

@@ Appendix C Limitations
Fig. 2 illustrates the current limitations of InterDyn on challenging scenarios. The first row shows that object consistency is not guaranteed, with the yellow toy changing shape when the tower falls over. The second row shows that when the object is not already in the frame, InterDyn sometimes generates an object that is not identifiable as belonging to a specific class. The third row shows a scenario in which it is particularly difficult to estimate the depth of the scene, causing InterDyn to drop the pen while it is already on top of the surface. Lastly, the fourth row shows a scenario in which a coin should be buried in flour, however, the object in our generation simply disappears.
Appendix D Video Class Analysis
As mentioned in Sec. 4.4, the Something-Something-v2 dataset includes videos of human-object interactions, such as “pushing [something] from left to right”. These phrases serve as class labels, enabling us to analyze the video classes where InterDyn (256$\times$384) performs best and worst, in terms of motion fidelity. Table 5 presents the 20 best and 20 worst-performing classes based on motion fidelity, alongside the number of videos for each class in the validation set and the average motion fidelity score.
First of all, we notice that InterDyn generally performs less effectively on underrepresented classes within the dataset, while at the same time, many of these underrepresented classes involve complex dynamics, such as spinning, burying, or folding objects. From Tab. 5, we observe that InterDyn (256$\times$384) performs best on simpler motions, particularly translations relative to the camera. This includes actions like moving an object up, down, or across, as well as dropping objects, moving objects toward or away from the camera, and picking up objects. Conversely, InterDyn struggles with classes involving highly complex dynamics, such as non-translational object-object interactions (e.g., throwing one object at another, burying an object, or poking structures made up of multiple components like stacks or toy towers). It also underperforms on spinning objects or their parts (e.g. a fidget spinner or a ceiling fan). We believe this is not an inherent limitation of InterDyn but rather the result of a combination of factors: the dataset’s frame rate (FPS), the presence of motion blur, and the under-representation of these more unique classes.
Appendix E Ethical Considerations
Video generation models, such as InterDyn, could be misused for creating deepfake videos that spread misinformation or compromise trust. Such applications pose a large risk for society and highlight the importance of developing ethical guidelines that can serve as safeguards when developing video generation models as well as when applying them to downstream tasks. On the positive side, advances such as InterDyn could be used to enhance storytelling in entertainment or aid in deepfake detection by identifying motion-dynamic inconsistencies. By advancing our understanding of interactive dynamics responsibly, we can amplify their positive societal impact while mitigating risks.
[…]
Order Label Count Avg motion fidel.
1 Moving something down 182 0.86
2 Pulling something from right to left 57 0.84
3 Moving something up 197 0.82
4 Pulling something from left to right 83 0.82
5 Holding something over something 165 0.80
6 Holding something 103 0.80
7 Moving something across a surface without it falling down 26 0.79
8 Pushing something from left to right 123 0.79
9 Holding something in front of something 138 0.78
10 Pushing something from right to left 122 0.77
11 Putting something on a surface 85 0.77
12 Moving something across a surface until it falls down 28 0.77
13 Lifting something with something on it 369 0.77
14 Squeezing something 216 0.77
15 Lifting something up completely without letting it drop down 66 0.75
16 Throwing something in the air and letting it fall 6 0.75
17 Moving something closer to something 105 0.75
18 Holding something next to something 135 0.75
@@ Appendix E Ethical Considerations
19 Putting something that can’t roll onto a slanted surface, so it stays where it is 15 0.75
20 Trying to bend something unbendable so nothing happens 74 0.74
114 Spinning something so it continues spinning 51 0.47
115 Poking something so that it falls over 42 0.46
116 Pulling something out of something 33 0.46
117 Folding something 187 0.46
118 Poking something so it slightly moves 71 0.45
119 Spinning something that quickly stops spinning 47 0.45
120 Taking something out of something 66 0.45
121 Unfolding something 122 0.44
122 Putting something, something, and something on the table 60 0.44
123 Piling something up 27 0.43
124 Something being deflected from something 10 0.41
125 Poking something so lightly that it doesn’t or almost doesn’t move 83 0.41
126 Burying something in something 4 0.41
127 Showing something next to something 19 0.40
128 Pushing something so it spins 23 0.39
129 Poking something so that it spins around 7 0.39
130 Putting number of something onto something 5 0.37
131 Poking a stack of something so the stack collapses 8 0.34
132 Showing something on top of something 14 0.34
133 Wiping something off of something 9 0.32
@@
Table 5: Motion fidelity for different action classes on the Something-Something-v2 dataset. The table shows the top and bottom 25 categories in terms of motion fidelity, together with the number of samples for that category in the validation set. The top categories contain many translation dynamics with respect to the camera, such as moving something up or from left to right. The bottom categories contain very complex dynamics such as spinning, burying, or showing an object from behind something.

### v3

@@ Appendix C Limitations
InterDyn performs best on translations relative to the camera; dropping objects, moving objects toward or away from the camera, and picking up objects. InterDyn struggles with complex non-translational interactions (e.g. throwing one object at another, burying an object, or poking a tower of stacked objects). It also underperforms in scenarios where depth is ambiguous in the input image, see Fig. S2.
Label Count MF $\uparrow$ (avg.)
Moving something down 182 0.86
Pulling something from right to left 57 0.84
Moving something up 197 0.82
Pulling something from left to right 83 0.82
Holding something over something 165 0.80
Holding something 103 0.80
Moving something across a surface without it falling down 26 0.79
Pushing something from left to right 123 0.79
Holding something in front of something 138 0.78
Pushing something from right to left 122 0.77
Putting something on a surface 85 0.77
Moving something across a surface until it falls down 28 0.77
Lifting something with something on it 369 0.77
Squeezing something 216 0.77
Lifting something up completely without letting it drop down 66 0.75
Throwing something in the air and letting it fall 6 0.75
Moving something closer to something 105 0.75
Holding something next to something 135 0.75
@@
Putting something that can’t roll onto a slanted surface, so it stays where it is 15 0.75
Trying to bend something unbendable so nothing happens 74 0.74
(a) Top 20 categories. Contains many translation dynamics with respect to the camera, such as moving something up or from left to right.
Label Count MF $\uparrow$ (avg.)
Spinning something so it continues spinning 51 0.47
Poking something so that it falls over 42 0.46
Pulling something out of something 33 0.46
Folding something 187 0.46
Poking something so it slightly moves 71 0.45
Spinning something that quickly stops spinning 47 0.45
Taking something out of something 66 0.45
Unfolding something 122 0.44
Putting something, something, and something on the table 60 0.44
Piling something up 27 0.43
Something being deflected from something 10 0.41
Poking something so lightly that it doesn’t or almost doesn’t move 83 0.41
Burying something in something 4 0.41
Showing something next to something 19 0.40
Pushing something so it spins 23 0.39
Poking something so that it spins around 7 0.39
Putting number of something onto something 5 0.37
Poking a stack of something so the stack collapses 8 0.34
Showing something on top of something 14 0.34
Wiping something off of something 9 0.32
(b) Bottom 20 categories. Contains very complex dynamics such as spinning, burying, or showing an object from behind something.
@@
Table S3: Motion fidelity for different action classes on the Something-Something-v2 dataset. The table shows the top and bottom 20 categories, together with the number of samples for that category in the validation set.
We report the 20 video classes on which InterDyn 256$\times$384 performs best and worst in terms of motion fidelity, alongside the number of videos for each class in the validation set and the average motion fidelity score for that class, see Tab. S3. We generally notice that InterDyn performs less effectively on underrepresented classes within the dataset, while at the same time, many of these underrepresented classes involve complex dynamics, such as spinning, burying, or folding objects.
Figure S3: Additional qualitative results on SSV2. We show multiple challenging examples, such as (from top to bottom): spinning a fidget spinner, tilting a sleek ridged object, squeezing a ball despite receiving an incomplete control signal, hand object-object interaction, zooming in, squeezing a sponge, dropping a hairband, or hand object-object interaction despite receiving a sparse control signal.
Appendix D Additional Results
We show additional qualitative examples in Fig. S3.
Appendix E Pretending class
Similar to our baseline CosHand [70], we removed the “pretending” class from SSV2 for training and validation, to avoid introducing ambiguous training signals. To compare its performance on this class to our results in Tab. 1, we run InterDyn on the “pretending” class in the validation split (828 samples for 256$\times$384 and 1156 for 256$\times$256), see Tab. S4. While the generations stay consistent in terms of image quality, we notice that motion fidelity is lower. Unfortunately, since FID, KID, FVD, and KVD compare distributions and are heavily dependent on the number of data samples, we cannot directly compare these metrics to those reported in Tab. 1.
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$ FVD $\downarrow$ KVD $\downarrow$ Motion Fidelity $\uparrow$ [93]
Seer [23] 0.357 9.42 0.657 74.86 0.060 640.06 147.07
DynamiCrafter [89] 34.96 0.016 314.05 34.22
CosHand-Independent [70] 0.620 16.79 0.310 9.85 0.003 123.59 15.97 0.396
CosHand-Autoregressive [70] 0.534 14.80 0.410 24.10 0.012 139.07 11.18 0.512
Ours 256$\times$256 0.666 18.52 0.256 14.29 0.004 49.02 -0.131 0.573
Ours 256$\times$384 0.683 18.99 0.249 17.39 0.004 64.18 -0.450 0.572
Table S4: Quantitative comparison on the “pretending” class of SSV2. We compare against Seer [23], DynamiCrafter [89] and two video extensions of our baseline CosHand [70]. Methods denoted with $\dagger$ do not use SSV2 as their training dataset.